The two Taylor girls were born in a consecutive, even years. If the product of their ages is 224, how old is each girl.
Answer by ankor@dixie-net.com(22740)

The two Taylor girls were born in a consecutive, even years.
If the product of their ages is 224, how old is each girl.
Let x = younger girl
then
(x+2) = older girl
:
x(x+2) = 224
:
x^2 + 2x - 224 = 0
Factors to
(x+14)(x-12) = 0
positive solution
x = 12 yrs is the younger girl
then, obviously:
14 yrs is the older